OpenStudy (camerondoherty):

Four polynomial functions are shown below: \(f(x) = x^3 + 3x^2 - 4x - 12\) \(g(x) = x^3 - 3x - 2\) \(p(x) = x^2 + 3x + 2\) \(t(x) = x^2 + x - 2\) Which of the following functions has -1 and -2 as its zeros?

OpenStudy (camerondoherty):

I'm thinking either C or D

OpenStudy (bibby):

not sure how to factor the first 2 but p(x) factors to (x+2)(x+1)

OpenStudy (camerondoherty):

Its not the last one

OpenStudy (bibby):

\(p(x) = x^2 + 3x + 2\) \(x^2 + 3x + 2=0\) \((x+2)(x+1)=0\) x={-1,-2}

OpenStudy (camerondoherty):

Oh, thanks lol

OpenStudy (bibby):

np np :)